Circular No. 5899

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 H.-U. Zimmermann, B. Aschenbach, G. Hasinger, W. Pietsch, P. Pre-

dehl, and J. Trümper, Max-Planck-Institut für Extraterrestrische Physik; W. Lewin and E. Magnier, Massachusetts Institute of Technology; J. van Paradijs, University of Amsterdam; G. Fabbiano, Center for Astrophysics; L. Lubin, Princeton University; and R. Petre, Goddard Space Flight Center, report: “On Nov. 1 and 2, SN 1993J was again observed by the x-ray satellite ROSAT. Since the last observation in May, the PSPC count rate in the energy band 0.1-2.4 keV has decreased from 0.053 ± 0.0025 to 0.035 ± 0.0014 count/s. The energy spectrum has become significantly softer (with kT ∼ 0.5 keV for a thermal bremsstrahlung model, compared to > 7 keV for the May measurements). The low-energy x-ray absorption corresponds to a neutral hydrogen column density of ∼ 3 × 10^{21} cm^{−2}, which is about a factor of 6 higher than in the May observation. The change in the absorbing column density and the low temperature seem to indicate that part of the wind material, lost by the progenitor star and highly ionized by the radiation from the initial shock, has started to recombine.”

PERIODIC COMET ENCKE

Ephemeris from orbital elements on MPC 18256:

1993/94 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 29 22^{h}40.^{m}37 + 7^{o}28.^{′}4 0.953 1.457 97.^{o}4 42.^{o}2 12.1
Dec. 4 22 36.29 + 6 36.3 0.956 1.388 91.2  45.2 11.7
9 22 33.33 + 5 50.8 0.956 1.316 85.4  48.2 11.3
14 22 31.34 + 5 11.5 0.954 1.242 79.7  51.2 10.9
19 22 30.18 + 4 37.6 0.947 1.166 74.2  54.3 10.5
24 22 29.64 + 4 07.8 0.936 1.087 69.0  57.6 10.1
29 22 29.48 + 3 40.0 0.918 1.006 63.8  61.3  9.6
Jan. 3 22 29.41 + 3 11.6 0.893 0.921 58.6  65.6  9.2
8 22 28.98 + 2 38.5 0.861 0.834 53.2  70.9  8.8
13 22 27.57 + 1 53.9 0.821 0.744 47.6  77.7  8.3
18 22 24.12 + 0 46.6 0.774 0.651 41.4  86.9  7.9
23 22 16.84 − 1 03.1 0.721 0.558 33.9  99.9  7.4
28 22 02.96 − 4 07.0 0.669 0.467 24.5 119.2  7.1

1993 November 30 Daniel W. E. Green